the measure of the largest angle of a triangle is nine times the measure of the smallest angle. The measure of the third angle is equal to the difference of the largest and the smallest what are the measures of the angles
Answer by mananth(16949)

let smallest angle be x
largest angle = 9x
Third angle = 9x-x=8x
...
x+9x+8x=180
18x=180
/18
x=180/18
x= 10
the angles are 10,80,90